Please show work.

Company A:
Demand is 80-.2P=Q
Fixed Cost is $200
Variable Cost is $60

find the following:

inverse demand function
total revenue function
marginal revenue function
revenue max quantity
revenue max price
total cost function
marginal cost
profit max quantity
profit max price
total revenue
total cost
profit or loss
Answer #1

Q = 80 -0.2P 6.2p= 80 - Q p= 6oo -5Q Inverse demand : p=400 - 59 Potal Revenue : PQ = 4000 - 5a Marginal Revenue MR= d (PQ) =